Shift Checklist HowTos
From SBS wiki
<<SBS Main<<Shift Takers (Shift Leader)
This checklist guide is intended to be a help in filling out the checklist, but don't take its instructions too literally.
- Your name: check your JLab ID to make sure you have spelled your name correctly. By the end of and Owl shift even this can be difficult.
- Shift: I hope the answer to this is obvious. If not please call the guard shack and request to be escorted off site.
- Shift Leader: your name should go here. If not, rebuke the shift leader for obvious neglect of duty and inquire as to why the SL is not performing the shift checklist per protocol.
- Target operator: the name of the target operator goes here. Check the TO's ID badge to ensure that the name and picture are a match and that we don't have outside spies trying to steal our data. If the credentials do not match or if the ID badge resembles that in the figure below, notify the RC, the crew chief and all the collaboration.
- Beam energy, current, positions (BPM1H04A/E X/Y) can be seen in Figure 2. This information is also available on the "BEAMLINE" section of the Hall A General Tools GUI. Extra credit will be given if the beam energy is converted to calories, the beam current to the number of elementary charges per second and the positions to units of Planck length.
- Energy lock: to find the energy lock screen type "energy lock" into the jmenu search bar and select "Energy Locks". The screen should look as pictured on the left below with the 1A and 2A green lights on if the ARC l and 2 energy locks are engaged. If you have trouble finding the energy lock, then other hair products will work as well.
- Target vacuum: ask the target operator. It is available on the target screen and should be in the 10^-7 Torr or less. If the TO doesn't know where to find it, send them home and call HR to have their paycheck stopped immediately.
- Fast raster size at target: due to the quadrupole optics of the Hall A beamline downstream of the raster, the MCC setpoint for the raster seldom reflects the actual size at the target. The actual size is communicated via the whiteboard or by word of mouth but should usually be 2mm x 2mm. If you aren't sure, ask someone who knows (in this situation praying doesn't count as "asking someone who knows").
- Fast raster MCC units: this is available on the Hall A status screen (see the yellow outline in Figure 3.) Psst, this is what MCC thinks the size is. Don't give away our secret.
- Target alarm handler: ask the TO if the target alarm handler is on, visible and working. If the TO is asleep, gently turn the alarm volume to maximum and press the "test sound" button.
- Alarm Handler status: the Hall A general alarm handler should be active on hadesk9. Check to make sure it is there, active and that the volume is up by pressing the "test sound" drop down menu option. No humor here. What is wrong with you? Are you expecting humor on every step of the list?
- Target in use (ask TO). Ask the TO is right in the question...so who needs guidance on this one. Good answers from the TO include optics, carbon 2mm hole, carbon 5mm hole, LH2, LD2, LH2+rad, LD2+rad and dummy. If the TO answers "dummy", don't take it as an insult; write it down.
- Target temperature/pressure: these are available on the target control GUI. In case you haven't been for a checkup lately, your target temperature and pressure should be 98.6°F and 120/80 mmHg. We want a healthy shift crew!
- The target oven temperature and magnetic field strength and direction should be read from the strip charts on hadesk7. There is a log entry here with the relevant EPICS variables: https://logbooks.jlab.org/entry/4115796
- Hall A VESDA system OK? This panel is located on the wall as you come into the Hall A Counting House front room and should have all lights green and in the lowest alert setting as shown in the figure below. VESDA stands for very early smoke detection alarm, so no trying to sneak a smoke inside the LHRS hut (even if Clint Eastwood is cool with it).
- DAQ configuration: this is selected during Prestart of the DAQ and can be read from the halog start and end of run logs. It will generally be GEnII-NoSBSGems during production. See example log entry below with configuration circled. The configuration can also be read directly off the CODA control GUI as seen circled in the orange ellipse in Figure-6. Time to relax and smile without a humorous stimulation. Try it yourself: pull both edges of your mouth upward. Didn't work? Ok, then, try this.
- Trigger configuration: this is set by the prescales which can be obtained by looking at the latest EPICS start of run halog and copy/pasting the line with the prescales near the top as shown in the figure below. These can also be obtained via the prescale GUI on the DAQ computer. You are getting close to the end. I know you can do it! Keep going. The collaboration is counting on you! Translation: Don't be known as the scoundrel that ruined GEn.
- BigBite current, polarity: this can be read directly off the BigBite controls GUI which should be open on hadesk8. If it is not, you can open it using the instructions here. Note that for BigBite, the current magnitude and sign (polarity) are given separately, whereas for SBS, you get it from the sign of the readback current.
- SBS current, polarity: this can be read directly off the SBS controls GUI which should be open on hadesk8. If it is not, you can open it using the instructions here. Note that for BigBite, the current magnitude and sign (polarity) are given separately, whereas for SBS, you get it from the sign of the readback current. Speaking of current,
- HV checked and looks good: go to the HV GUI open on hadesk9 and go through the various tabs to check the HV for the BBCal shower and preshower, the INFN and UVA GEMs, the GRINCH and HCal. Usually, these are all on. Leave a comment if they are not with a reason why. If the HV GUI is not open instructions for opening it are here.
- HCal and BBCal thresholds are both found on the threshold GUI open on hadesk9. Check the readback in mV for "HCal Overlapping Region Threshold" and "BBCalHi Discriminator 1(2)" values. If the GUI is not open, instructions for opening it can be found here.
- TS1-BB and L1A scaler rates: read directly off the scaler GUI on the "BCM" tab. This should be open on the overhead TVs (hatv1). If the scaler GUI is not open, instructions for opening it are here.
- Scaler rates reasonable? Wow! what a subjective question. Guidelines: the scalar rates are reasonable if the TS1-BB rates are <5 kHz and the L1A is close to the TS1-BB rate. If L1A is significantly lower than TS1-BB, then the livetime is likely low. If the ratio (L1A/TS1-BB)>=~90%, then L1A scaler is also reasonable.
- GEMs need proper gas mixtures flowing through them. On any a-onl@aonlX machine run gosbs and CheckGEMGas.sh. The terminal will then tell you if everything is ok or not.
- The parity DAQ should be running on vnc. Login as adaq@adaq3 then type "vncviewer -Shared :40". CODA should be running in the vnc. If the parity DAQ is not running, contact the parity DAQ subsystem expert.